# Model-Based Analysis of Mining Fairness in a Blockchain

### v2

@@ 4.1 Examining the Scale of Forks
The model-based calculation disregards the impact of large-scale forks. Specifically, it assumes that the number of blocks per round is at most two. In this study, we investigate the effects of large-scale forks.
Regarding the Scale of Forks First, we establish some facts regarding fork rates. Let the hashrate of miner $i$, where $i\in V$, be $M_{i}$. Let the total network hashrate be $M_{all}$. Additionally, let the probability of successfully generating a block with one hash calculation be $p$. Then, the average number of hash calculations required to generate a block is $1/p$. Therefore, the following equation holds:
@@
where $P(C_{i}=1)$ denotes the probability that no forks occur. Then, the probability $P(C_{i}\neq 1)$ that a fork occurs is
@@
The probability that the number of blocks in a round will be 3 or more satisfies the following inequality:
@@
$\frac{T_{W,i}}{T}\ \text{is constant}]{p\ \to\ 0}1-(1+\frac{T_{W,i}}{T})e^{-\frac{T_{W,i}}{T}}.$ (28)
When the number of blocks in a round is 2 or more, at least 3 hash calculations succeed before all the blocks are fully shared, thus satisfying (22). Equation (27) is obtained by substituting (17) into (26).
From (28), it follows that the probability that the number of blocks in a round will be 2 satisfies the following inequality:
@@
In the model-based calculation of mining fairness, we consider cases in which the number of blocks per round is two or fewer, thereby ignoring $I_{3}$. Thus, we compare $I_{3}$ with $I_{1}$ and $I_{2}$. Figs. 3 and 4 show $I_{3}/(I_{1}+I_{2})$ and $I_{3}/I_{2}$, respectively, as $d/T$ is varied from 00 to $1$. The specific numerical values are listed in Table I. It is evident that as $d/T$ decreases, the influence of $I_{3}$ diminishes. Although it is not directly demonstrated herein whether the model-based calculation can determine actual mining fairness, it can be inferred that the model-based calculation will be effective in scenarios where $d/T$ is small.
4.2 Network Composed of Two Miners
@@ 4.2.1 Model-Based Calculation
Afterward, we calculate the probability $W_{AB}$ that the block generated by $Miner_{A}$ is incorporated into the main chain if $Miner_{B}$ generates a block by forking immediately after $Miner_{A}$ starts a round. For simplicity, we assume that the block of $Miner_{B}$ conflicts with the chain of $Miner_{A}$. That is, we ignore the case in which the block height of $Miner_{B}$ is smaller than that of $Miner_{A}$.
[…]
@@ 4.2.2 Simulation Settings
Based on the blockchain network simulator SimBlock [19], we developed another event-driven simulator composed of two miners. Our simulator is capable of simulating forks of any scale, similar to those in real blockchain systems.
@@
We also examined variations in $\alpha_{A}$, i.e., the proportion of hashrate of miner $A$, with values of $0.1$ and $0.3$. A value of $0.5$ was not considered because, in this case, mining fairness is completely maintained owing to the symmetry of the network.
Each simulation consisted of 10 billion rounds.
TABLE II: Errors Between Simulation and Model-Based Calculation Results for $LF_{1}$.
@@ 4.2.3 Validation Results
Based on the error values, it can be seen that the model-based calculation can compute mining fairness with high accuracy. It can also be observed that the accuracy deteriorates as $d/T$ increases. This is likely because, as seen in Section 4.1, the impact of having more than three blocks per round becomes more significant as $d/T$ increases.
4.3 Network Composed of 10 Miners
In this section, we validate the model-based calculation of mining fairness on a network composed of 10 miners. Compared to a network with two miners, a network with ten miners introduces additional elements such as tie-breaking rules, hashrate distribution, and block propagation time. This allows us to demonstrate that the proposed model-based calculation method is effective even in more complex networks.
@@ 4.3.1 $W_{ij}$ in a Network Composed of Multiple Miners
In a network with more than two miners, it is necessary to consider tie-breaking rules. Here, we demonstrate how to determine $W_{ij}$ for a network with multiple miners according to different tie-breaking rules. It is assumed that all forks cause chain ties.
@@
4.3.2 Simulation Settings
The simulator used in this validation was an extended version of a network simulator composed of two miners. The number of miners was set to 10. The hashrate distribution was based on that of Bitcoin [20]. The hashrate distribution settings are shown in Fig. 5.
Each simulation consisted of 10 billion rounds.
@@ 4.3.2 Simulation Settings
Exiting method 1.1644 1.16133 1.15597 1.15099
@@ 4.3.3 Validation Results
We conducted 50 simulation experiments for each validation target. The errors between the simulation and model-based calculation results are shown in Tables III and IV. Here, error is defined as the relative error as in (53). The difference from the previous validation is that the number of elements in the $LF$ vector is changed from two to ten, and we investigate not only the $LF$ vector but also the round-start-rate vector.
@@
First, we examine the round start rate. It can be seen that the model-based calculations match the simulation results with high accuracy under all conditions.
@@
Next, we examine mining fairness. From the tables, it can be seen that the model-based calculation method is able to compute mining fairness with high accuracy. It can also be observed that the accuracy deteriorates as $d/T$ increases.
@@
Furthermore, it is found that the accuracy of the mining fairness calculations is not as high as that of the round-start-rate calculations or of the mining fairness calculations for a network composed of two miners. This is because the mining fairness calculation requires $W_{ij}$.
@@
Additionally, it is observed that the first-seen rule is more accurate than the random rule or the last-generated rule. This finding indicates that the calculation of $W_{ij}$ based on the first-seen rule is superior. This is because the influence of blocks up to the second one in a given round is stronger under the first-seen rule.

### v3

@@ 5.1 Examining the Scale of Forks
The model-based calculation disregards the impact of large-scale forks. In particular, it assumes that the number of blocks per round is at most two. In this study, we investigate the effects of large-scale forks.
Regarding the Scale of Forks: First, we establish some facts regarding fork rates. Let the hashrate of miner $i$, where $i\in V$, be $M_{i}$. Let the total network hashrate be $M_{all}$. Additionally, let the probability of successfully generating a block with one hash calculation be $p$. Then, the average number of hash calculations required to generate a block is $1/p$. Therefore, the following equation holds:
@@
where $P(C_{i}=1)$ denotes the probability that no forks occur. Then, the probability $P(C_{i}\neq 1)$ that a fork occurs is as follows:
@@
The probability that the number of blocks in a round will be three or more satisfies the following inequality:
@@
$\xrightarrow[\frac{T_{W,i}}{T}\ \text{is constant}]{p\ \to\ 0}1-(1+\frac{T_{W,i}}{T})e^{-\frac{T_{W,i}}{T}}.$ (28)
When the number of blocks in a round is two or more, at least three hash calculations succeed before all the blocks are fully shared, thus satisfying (22). Equation (27) is obtained by substituting (17) into (26).
From (28), it follows that the probability that the number of blocks in a round will be two satisfies the following inequality:
@@
In the model-based calculation of mining fairness, we consider cases in which the number of blocks per round is two or fewer, thereby ignoring $I_{3}$. Thus, we compare $I_{3}$ with $I_{1}$ and $I_{2}$. Figs. 3 and 4 show $I_{3}/(I_{1}+I_{2})$ and $I_{3}/I_{2}$, respectively, as $d/T$ is varied from 0 to $1$. The specific numerical values are listed in Table I. It is evident that as $d/T$ decreases, the influence of $I_{3}$ diminishes. Although it is not directly demonstrated herein whether the model-based calculation can determine actual mining fairness, it can be inferred that the model-based calculation will be effective in scenarios where $d/T$ is small.
5.2 Networks Composed of Two Miners
@@ 5.2.1 Model-Based Calculation
Thereafter, we calculate the probability $W_{AB}$ that the block generated by $Miner_{A}$ is incorporated into the main chain if $Miner_{B}$ generates a block by forking immediately after $Miner_{A}$ starts a round. For simplicity, we assume that the block of $Miner_{B}$ conflicts with the chain of $Miner_{A}$; that is, we ignore the case in which the block height of $Miner_{B}$ is smaller than that of $Miner_{A}$.
$W_{AB}=$ $\overset{1}{\alpha_{A}}\overset{2}{\alpha_{A}}+\overset{1}{\alpha_{A}}\overset{2}{\alpha_{B}}(1-f)$
$+\overset{1}{\alpha_{A}}\overset{1}{\alpha_{B}}f\{\overset{2}{\alpha_{A}}\overset{3}{\alpha_{A}}+\overset{2}{\alpha_{A}}\overset{3}{\alpha_{B}}(1-f)$
$+\overset{2}{\alpha_{A}}\overset{2}{\alpha_{B}}f(\cdots)+\overset{2}{\alpha_{B}}\overset{2}{\alpha_{A}}f(\cdots)\}$
$+\overset{1}{\alpha_{B}}\overset{1}{\alpha_{A}}f\{\overset{2}{\alpha_{A}}\overset{3}{\alpha_{A}}+\overset{2}{\alpha_{A}}\overset{3}{\alpha_{B}}(1-f)$
$+\overset{2}{\alpha_{A}}\overset{2}{\alpha_{B}}f(\cdots)+\overset{2}{\alpha_{B}}\overset{2}{\alpha_{A}}f(\cdots)\}$ (45)
@@ 5.2.2 Simulation Settings
Based on the blockchain network simulator SimBlock [25], we developed another event-driven simulator composed of two miners. Our simulator can simulate forks of any scale, similar to those in real blockchain systems.
@@
We also examined variations in $\alpha_{A}$, i.e., the proportion of hashrate of miner $A$, with values of $0.1$ and $0.3$. A value of $0.5$ was not considered because, in this case, mining fairness is completely maintained because of the symmetry of the network.
Each simulation consisted of ten billion rounds.
TABLE II: Errors Between Simulation and Model-Based Calculation Results of $LF_{1}$ for Network of Two Miners.
@@ 5.2.3 Validation Results
The error values indicate that the model-based calculation can compute mining fairness with high accuracy. Furthermore, it can be observed that the accuracy deteriorates as $d/T$ increases. This is likely because, as seen in Section 5.1, the impact of having more than three blocks per round becomes more significant as $d/T$ increases.
5.3 Network Composed of Ten Miners
In this section, we validate the model-based calculation of mining fairness on a network comprising ten miners. Compared to a network with two miners, a network with ten miners introduces additional elements, including tie-breaking rules, hashrate distribution, and block propagation time; this allows us to demonstrate that the proposed model-based calculation method is effective even in more complex networks. Furthermore, we compare our method against a state-of-the-art approach [14], highlighting its advantageous performance.
@@ 5.3.1 $W_{ij}$ in a Network Composed of Multiple Miners
In a network with more than two miners, it is necessary to consider tie-breaking rules. Here, we demonstrate how to determine $W_{ij}$ for a network with multiple miners according to different tie-breaking rules. We assume that all forks cause chain ties.
@@
Existing method 1.1644 1.16133 1.15597 1.15099
5.3.2 Simulation Settings
The simulator used in this validation was an extended version of a network simulator composed of two miners. The number of miners was set to ten. The hashrate distribution was based on that of Bitcoin [26]. The hashrate distribution settings are illustrated in Fig. 5.
Each simulation consisted of ten billion rounds.
@@ 5.3.3 Validation Results
We conducted 50 simulation experiments for each validation target. The errors between the simulation and model-based calculation results are listed in Tables III and IV. Here, error is defined as the relative error as in (53). The difference from the previous validation is that the number of elements in the $LF$ vector is changed from two to ten, and we investigate not only the $LF$ vector but also the round-start-rate vector.
@@
First, we examine the round start rate. As observed, the model-based calculations match the simulation results with high accuracy under all conditions.
@@
Next, we examine mining fairness. The tables demonstrate that the model-based calculation method can compute mining fairness with high accuracy; additionally, it can be observed that the accuracy deteriorates as $d/T$ increases.
@@
Furthermore, it is observed that the accuracy of the mining fairness calculations is not as high as that of the round-start-rate calculations or of the mining fairness calculations for a network composed of two miners; this is because the mining fairness calculation requires $W_{ij}$.
@@
Additionally, it is observed that the first-seen rule is more accurate than the random rule or the last-generated rule. This finding indicates that the calculation of $W_{ij}$ based on the first-seen rule is superior; this is because the influence of blocks up to the second one in a given round is stronger under the first-seen rule.